Formic acid, HCHO2 (46.02 g / mole) is a convenient source of small quantities of carbon monoxide. When warmed with sulfuric acid, formic acid decomposes to give Co (28.01 g / mole) gas. HCHO2 () - H20 () + CO (g) If 5.25 L of carbon monoxide was collected over water at 25 ° C and 689 mmHg, how many grams of formic acid were consumed? The vapor pressure of water at 23 ° C is 23.8 mmHg.
1
Expert's answer
2020-06-03T13:34:21-0400

p(CO)=p(total)-p(H2O)=689-23.8 =665,2.

pV=nRT.

n=pV/RT= 665.2 x 5250 / 62400 x 298 = 0,187806531 moles of CO.

0,187806531 x 46.02 = 8,64285656 g.